# Fragment-Masked Molecular Optimization

### v2

@@ Introduction
To explore whether the molecules generated or optimized using different methods achieved a certain confidence level instead of merely evaluating the methods based on numerical values, we visually analyzed the molecular structures generated using our optimization method and baselines across different cell lines. As shown in Fig. 3, our method generates unique molecules for each cell line, ensuring that the optimization process does not converge to the same local optimum across different cell lines. In comparison, the molecules generated using other methods were generally similar, highlighting that our method is able to optimize molecular structures based on the specific response values of each cell line, thereby achieving more effective and suitable molecular structures. Furthermore, our method’s predicted IC50 values remain consistently low across different cell lines, indicating that our optimized molecules have a competitive advantage compared to de novo designed molecules. For example, for cell line 906792, our method generated a molecule with an IC50 value of 0.48, while the other approaches generated different molecules with IC50 values ranging from 0.49 to 0.51.
@@
Since measuring the IC50 for all virtually generated molecules on 945 specific cell lines in a short time is impractical, we utilized the CLDR method to predict these values. To validate our method’s effectiveness, as shown in Fig. 4, we predicted the IC50 values for the cell lines 906792, 688006, and 949090, and visualized the mean and variance by assuming a Gaussian distribution. As a result, our proposed FMOP method demonstrated a strong competitiveness in the lower IC50 range. For example, in Fig. 4(a), the IC50 values for the molecules generated by FMOP fall between 0.47 and 0.53, while those generated using other methods mostly fall between 0.49 and 0.52. Additionally, the IC50 distribution of the original molecules (denoted as ’Origin’) ranged between 0.51 and 0.52. After optimization using our method, the mean IC50 of the molecules showed a significant improvement, approximately by 3%. The results generated by MOOD and DruM2D are represented by $\bullet$, indicating that only one single effective molecule was optimized for cell line 906792, resulting in a variance of 0.
Conclusion
@@
To address the PDD challenge of generating molecules, which requires screening a vast number of possible molecular structures, we proposed the FMOP method. To the best of our knowledge, the FMOP is the first optimization method for the PDD task. FMOP employs a regression-free diffusion model to conditionally sample the masked regions of molecules for optimization, effectively generating new molecules with similar scaffolds and improved efficacy. We optimized the molecules for all 945 cell lines on the GDSCv2. The overall experiments demonstrated that the in-silico optimization success rate reaches 94.4%, with an average efficacy increase of 5.3%. Additionally, we conducted extensive ablation studies and visualization experiments, proving that FMOP is an effective and robust molecular optimization method.

### v3

@@ IV-E Visualization Analysis
To explore whether the molecules generated or optimized using different methods achieved a certain confidence level instead of merely evaluating the methods based on numerical values, we visually analyzed the molecular structures generated using our optimization method and baselines across different cell lines. As shown in Fig. 5, our method generates unique molecules for each cell line, ensuring that the optimization process does not converge to the same local optimum across different cell lines. In comparison, the molecules generated using other methods were generally similar, highlighting that our method is able to optimize molecular structures based on the specific response values of each cell line, thereby achieving more effective and suitable molecular structures. Furthermore, our method’s predicted IC50 values remain consistently low across different cell lines, indicating that our optimized molecules have a competitive advantage compared to de novo designed molecules.
@@
Since measuring the IC50 for all virtually generated molecules on 985 specific cell lines in a short time is impractical, we utilized the CLDR method to predict these values. To validate our method’s effectiveness, as shown in Fig. 4 and Fig. 3, we predicted the IC50 values for the cell lines 906792, 687800, 684055 and 908149, then visualized the mean and variance by assuming a Gaussian distribution. As a result, our proposed FMOP method demonstrated a strong competitiveness in the lower IC50 range. For example, as shown in Fig. 4(c), the IC50 values for the molecules generated by FMOP range from 0.46 to 0.52, whereas the molecules generated by other methods typically include only one successfully optimized molecule (represented by $\bullet$). Additionally, the IC50 distribution of the original molecules (denoted as ’Origin’) ranged between 0.51 and 0.52.
Figure 8: Detailed interface for steps 6 and 7, where users interactively select specific atomic positions within a molecule for fragment optimization. The selected fragment must contain at least one atom, and the optimization region is not limited to specific fragment libraries or functional groups.
IV-F Case Study
We have deployed our method on the Beishenglai platform  [^1] The online platform can be accessed at https://www.baishenglai.com., a drug discovery platform based on deep learning models. The platform supports key drug discovery processes: generation, optimization, prediction, and retrosynthesis. To showcase the practical application of our method, we present a case study. In this case study, our FMOP method was employed to optimize the efficacy of the molecule Z-LLNle-CHO (Compound ID: 16760646, uniquely identified in the PubChem database)  [^2] The SMILES representation of Z-LLNle-CHO is CCCCC(C=O)NC(=O) [C@H](CC(C)C)NC(=O)[C@@H](CC(C)C)NC(=O)OCC1=CC=CC=C1. for a specific cancer type, Mo-T (ID: 908149 in the GDSCv2), while preserving the similarity of its molecular scaffold structure. The complete interface on the Beishenglai platform is shown in Fig. 7, where users input their target conditions and submit the optimization task.
Figure 9: Visualization of case study results, with normalized IC50 values annotated below each molecule.
The optimization process involves selecting the cell line, specifying the target IC50, choosing the number of molecules to generate, setting the diffusion time step, defining the guidance strength, and providing the original molecule. The number of molecules should be chosen based on the desired diversity, as a larger quantity may increase processing time. The time step controls optimization detail, with smaller values requiring more time. The guidance strength determines the optimization focus; higher values lead to more focused outputs but reduce diversity.
Fig. 8 illustrates the details for steps 6 and 7, allowing users to select atomic positions for fragment-masked optimization. The mask must include at least one atom, and the optimization region is not restricted to specific fragment libraries.
After task submission, the FMOP method was applied, and the results are shown in Fig. 9. Among the 100 optimized molecules,18 had an IC50 superior to the original, with the best-performing molecule (3) achieving an IC50 of 0.4721, an 11.4% improvement in efficacy. These results demonstrate the FMOP method’s ability to enhance key molecular properties and offer a flexible framework for optimization.
V Conclusion
@@
To address the PDD challenge of molecular optimization, which requires screening a vast number of possible molecular structures, we proposed the FMOP method. To the best of our knowledge, the FMOP is the first optimization method for the PDD task. FMOP employs a regression-free diffusion model to conditionally sample the masked regions of molecules for optimization, effectively generating new molecules with similar scaffolds and improved efficacy. We optimized the molecules for all 985 cell lines on the GDSCv2. The overall experiments demonstrated that the in-silico optimization success rate reaches 95.4%, with an average efficacy increase of 7.5%. Additionally, we conducted extensive ablation studies and visualization experiments, proving that FMOP is an effective and robust molecular optimization method.
Although the FMOP method has demonstrated exceptional performance in enhancing molecular efficacy, enabling optimization across multiple task conditions with a single training session, it still has limitations. First, our method is primarily suited for optimizing molecules with existing efficacy, assuming the target molecule exhibits some initial activity. In cases where efficacy is unclear or absent, the applicability of FMOP is limited. Second, FMOP relies on fragment masking for localized optimization, which may neglect the global structural requirements of the molecule.